### v3

# Quantum Adaptive Self-Attention for Quantum Transformer Models

Abstract
[…]
keywords:
quantum machine learning, quantum transformer, self-attention, time-series forecasting, NISQ, hybrid quantum-classical
articletype: Paper
1 Introduction
Transformer architectures [39] have become the backbone of modern deep learning, powering state-of-the-art models across natural language processing, computer vision, and sequential prediction. Their core strength lies in the self-attention mechanism, which dynamically models long-range dependencies. However, this comes at a significant computational cost—quadratic in sequence length—motivating the search for alternative attention paradigms.
Quantum computing offers a compelling candidate: parameterized quantum circuits (PQCs) operate in exponentially large Hilbert spaces and can represent correlations that are classically intractable to express compactly [22]. Recent work has explored quantum-enhanced sequence models, including quantum LSTMs [15] and quantum Transformer variants [5], which replace classical components wholesale with variational quantum circuits. However, a fundamental question remains unanswered: how much quantum integration is actually beneficial, and where should it be placed?
In this work, we propose Quantum Adaptive Self-Attention (QASA), a hybrid architecture designed around a principle of architectural parsimony—the idea that maximal quantum benefit comes not from maximizing quantum resources, but from placing minimal quantum computation at the optimal position. QASA retains N1N{-}1 classical Transformer encoder layers and replaces only the value projection in the final layer with a PQC, using just 36 trainable quantum parameters. This stands in contrast to approaches such as QLSTM (128 quantum parameters) and QnnFormer (90 quantum parameters), which distribute quantum computation across the entire model.
Our systematic evaluation across nine synthetic benchmarks, one real-world dataset, two quantum baselines (QLSTM and QnnFormer) alongside a classical Transformer, and a comprehensive ablation study yields the following key contributions:
1.1 Contributions
[…]
  • Task-conditional quantum advantage taxonomy. We provide empirical evidence that quantum enhancement benefits specific signal classes—chaotic dynamics, noisy oscillations, and smooth trends—while classical Transformers remain superior for clean periodic signals and sharp discontinuities. This taxonomy offers practitioners a principled guide for when to deploy hybrid quantum-classical models.
  • Trainability and noise resilience of the parsimonious design. Gradient-variance analysis shows that adding a second entangling layer collapses the gradient magnitude by ${\sim}30\times$, approaching a barren plateau—so our single-layer choice is not merely sufficient but necessary for trainability. Under depolarizing noise at hardware-realistic levels ($p\approx 10^{-3}$$10^{-2}$), QASA’s performance remains stable or improves, indicating that the design is NISQ-compatible without requiring error correction.
  • Comprehensive empirical validation. QASA achieves the best MSE on 4 of 9 benchmarks (tied with the classical baseline), significantly outperforms QLSTM on the seasonal trend task (p=0.009p{=}0.009), and reduces MAE by 6.0% on the real-world ETTh1 dataset. These results hold across multiple random seeds and are supported by statistical significance tests, effect size analysis, and non-parametric Friedman tests.
Figure 1: Overview of the QASA architecture. The model consists of N1N{-}1 standard Transformer encoder layers followed by one Quantum Encoder Layer. The Quantum Encoder Layer replaces the value projection in self-attention with a parameterized quantum circuit (PQC) operating on 8 qubits with 4 variational layers, followed by a classical feedforward network with residual connections.
2 Related work
2.1 Deep Learning on Time-Series Data
Deep learning has significantly advanced the modeling and understanding of time-series data across various domains. Recurrent Neural Networks (RNNs) [33], particularly Long Short-Term Memory (LSTM) networks [23] and Gated Recurrent Units (GRUs) [18], have been widely adopted due to their ability to capture long-term temporal dependencies. However, their sequential nature limits parallelization and increases training time. To address this, temporal convolutional networks (TCNs) [27] and Transformer-based architectures [39] have demonstrated superior performance by enabling parallel processing and better long-range dependency modeling.
In industrial and scientific applications, Transformer variants such as the Temporal Fusion Transformer (TFT) [28] and Informer [40] have been applied to multi-horizon prediction tasks with irregular time intervals and exogenous variables. Similar time-series learning techniques have been applied successfully in molecular dynamics and open quantum systems [37, 38, 7].
Despite these advancements, challenges remain in modeling noisy, sparse, or irregularly-sampled sequences, motivating the development of hybrid models that integrate domain-specific priors with general-purpose deep learning architectures.
2.2 Quantum Deep Learning on Time-Series Data
While classical deep learning has achieved remarkable success in time-series analysis, it often requires extensive computational resources and large-scale data to generalize effectively. With the advent of quantum computing, Quantum Deep Learning (QDL) [3] has emerged as a promising paradigm for modeling complex temporal patterns, offering theoretical advantages in expressivity and optimization through quantum entanglement and superposition.
In the context of time-series data, several quantum-inspired architectures have been proposed. Variational Quantum Circuits (VQCs) [20] and Quantum Recurrent Neural Networks (QRNNs) [35] have been introduced to capture temporal dependencies using parameterized quantum gates and recurrent structures. A particularly notable development is the Quantum Long Short-Term Memory (QLSTM) network [15], which integrates quantum circuits into the gating mechanisms of classical LSTMs. Recent comprehensive benchmarking studies [19, 9] have compared VQC-based regression against classical GRU and LSTM models, finding that while classical models often lead in raw accuracy, VQCs exhibit higher robustness to noise and better generalization on small-scale datasets. Quantum-enhanced channel mixing has also been explored in state-space models for temporal forecasting [11], while quantum temporal convolutional networks have been applied to cross-sectional equity return prediction [14].
Recent work has further expanded quantum approaches for temporal data. Tang et al. [36] proposed QnnFormer, a multi-attention mechanism leveraging quantum circuits for long-term forecasting, achieving significant error reductions on benchmark datasets. Chakraborty and Heintz [6] integrated hybrid quantum-classical attention into patch-based transformers for multivariate forecasting. Chittoor et al. [17] developed QuLTSF for long-term forecasting with quantum neural networks, demonstrating advantages in convergence stability for non-stationary data. Laskar and Goel [26] investigated the impact of different entanglement topologies in shallow circuits for time-series prediction on NISQ hardware. In finance, quantum-enhanced architectures have been investigated for demand prediction and volatility modeling [32].
Beyond time series, quantum machine learning has been applied to sequential data in other domains, including EEG encoding [8, 10] and multimodal contrastive learning [13]. Chen and Kuo [12] demonstrated that Transformer-assisted learning can effectively capture Lindblad dynamics in open quantum systems, establishing a direct connection between Transformer architectures and quantum physics. Despite these advancements, research on Quantum Transformers for time-series data remains limited. Most existing studies focus on Quantum Visual Transformers (QViT) [16] or fully quantum architectures such as Quixer [25] in sequence modeling, leaving a gap in understanding how hybrid quantum-classical self-attention mechanisms can be practically deployed for temporal forecasting. Addressing this gap is the key focus of this work.
3 Methodology
Figure 2: The parameterized quantum circuit (PQC) used in QASA’s QuantumLayer. The circuit operates on n=8n{=}8 data qubits plus one ancilla qubit, with L=4L{=}4 variational layers. Each layer applies single-qubit rotations ($R_{X}$, $R_{Y}$, $R_{Z}$; see Eq. (19) for definitions) and CNOT entangling gates in a ring topology. Input features are encoded via $R_{X}$ and $R_{Z}$ rotations on each qubit, and the output is obtained by measuring $\langle Z_{i}\rangle$ on all data qubits.
Table 1: Quantum Adaptive Self-Attention (QASA) Transformer Architecture. Each row lists a layer with its operation and input/output tensor shapes. $L$ denotes the input sequence length (e.g., L=50L{=}50), $d$ is the hidden feature dimension (e.g., d=256d{=}256), and $n$ is the number of qubits in the quantum circuit (e.g., n=8n{=}8). ’Attn’ is an abbreviation for ’attention’.
Layer Operation Input Shape Output Shape
Input Raw sequence $(L,1)$ $(L,1)$
Linear Embedding Linear + LayerNorm $(L,1)$ $(L,d)$
Positional Encoding Add sinusoidal PE $(L,d)$ $(L,d)$
Transformer Layer $\times(N{-}1)$ Multihead Attn + FFN $(L,d)$ $(L,d)$
Quantum Encoder Layer
Self-Attn + QNN + FFN
$(L,d)$ $(L,d)$
   QuantumLayer (QNN)
Linear $\rightarrow\mathbb{R}^{n}$
PQC $\rightarrow\mathbb{R}^{n}$
Linear $\rightarrow\mathbb{R}^{d}$ + Residual
$(L,d)$ $(L,d)$
Final Linear Extract $h[L]$ and project $(d)$ $(1)$
3.1 Overview
We propose a hybrid quantum-classical Transformer model tailored for sequential data forecasting. The architecture is designed to capture temporal dependencies through classical attention mechanisms, while integrating a parameterized quantum circuit (PQC) to enhance the model’s expressiveness and representation power.
Given an input sequence $x\in\mathbb{R}^{L\times 1}$ of length $L$, our model predicts a scalar target $\hat{y}\in\mathbb{R}$ corresponding to the next value in the sequence.
3.2 Embedding and Positional Encoding
The input sequence is first projected into a high-dimensional space using a linear layer, followed by layer normalization:
$h_{0}=\text{LayerNorm}(W_{e}x+b_{e}),\quad h_{0}\in\mathbb{R}^{L\times d},$ (1)
where $d$ is the hidden dimension. We then apply fixed sinusoidal positional encoding to inject temporal order information:
$h_{0}\leftarrow h_{0}+PE,$ (2)
where $PE$ denotes the positional encoding matrix.
3.3 Transformer Encoder Layers
The encoder consists of $N$ layers, where the first N1N-1 layers are standard Transformer encoder layers defined as:
$h_{i}=\text{Transformer}(h_{i-1}),\quad i=1,\dots,N-1.$ (3)
The Transformer layer consists of two main sublayers: a multi-head self-attention mechanism and a position-wise feed-forward network (FFN). Each sublayer is wrapped with a residual connection and layer normalization. Formally, the computation of the $i$-th Transformer layer can be described as follows:
$z_{i}$ $=\text{LayerNorm}\left(h_{i-1}+\text{MultiHeadSelfAttention}(h_{i-1})\right),$ (4)
$h_{i}$ $=\text{LayerNorm}\left(z_{i}+\text{FFN}(z_{i})\right).$ (5)
The multi-head self-attention mechanism is defined as:
$\text{MultiHeadSelfAttention}(X)=\text{Concat}(head_{1},\dots,head_{H})W^{O},$ (6)
where each attention head is computed as:
$head_{j}=\text{Attention}(XW_{j}^{Q},XW_{j}^{K},XW_{j}^{V}),\quad j=1,\dots,H.$ (7)
In the self-attention mechanism, each input vector is linearly projected into three different spaces to form the query ($Q$), key ($K$), and value ($V$) matrices:
$Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V},$ (8)
where $X\in\mathbb{R}^{T\times d_{\text{model}}}$ is the input sequence (with $T$ tokens), and $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ are learnable projection matrices.
The roles of $Q$, $K$, and $V$ are as follows:
  • Query ($Q$): Represents the current token’s request for information.
  • Key ($K$): Represents the ”content” or identity of each token in the sequence.
  • Value ($V$): Represents the actual information to be retrieved.
The attention mechanism computes a similarity score between each query and all keys:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V.$ (9)
This allows the model to retrieve contextually relevant information by weighting each value according to the query-key similarity.
The feed-forward network (FFN) is applied to each position separately and identically:
$\text{FFN}(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2},$ (10)
where $W_{1},W_{2},b_{1},b_{2}$ are learnable parameters.
Thus, the Transformer block can be summarized as:
$z_{i}=\mathrm{LayerNorm}\big(h_{i-1}+\mathrm{MultiHeadSelfAttention}(h_{i-1})\big),$ (11)
$h_{i}=\mathrm{LayerNorm}\big(z_{i}+\mathrm{FFN}(z_{i})\big),$ (12)
$\text{Transformer}(h_{i-1})=h_{i}.$ (13)
Each Transformer encoder layer includes multi-head self-attention, residual connections, and feedforward networks with GELU activation.
3.4 Quantum Encoder Layer
The final encoder block is a quantum-enhanced encoder layer. It begins with a multi-head self-attention module:
$a=\text{MultiHeadAttention}(h_{N-1})=\text{Concat}(head_{1},\ldots,head_{h})W^{O},$ (14)
$\text{where }head_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$ (15)
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (16)
$h^{\prime}=\text{LayerNorm}(h_{N-1}+a).$ (17)
Each token vector $h^{\prime}_{i}\in\mathbb{R}^{d}$ is passed through a quantum layer. The quantum layer first linearly projects the input to a lower-dimensional space suitable for quantum encoding:
$h_{q}=\tanh(W_{q}h^{\prime}_{i}),$ (18)
where $h_{q}\in\mathbb{R}^{n}$ and $n$ is the number of qubits.
The vector $h_{q}$ is then passed as input to a parameterized quantum circuit (PQC), which is defined over $n+1$ qubits and $L_{q}$ layers of unitary operations. The PQC applies data re-uploading using $R_{X}$ and $R_{Z}$ gates, and introduces entanglement using CNOT and $R_{Y}$ gates. Throughout this paper, the single-qubit rotation gates are defined as
$R_{X}(\theta)=e^{-i\theta X/2},\quad R_{Y}(\theta)=e^{-i\theta Y/2},\quad R_{Z}(\theta)=e^{-i\theta Z/2},$ (19)
where $X,Y,Z$ are the Pauli matrices. The final output is computed as the expectation values of Pauli-$Z$ operators on the first $n$ qubits:
$\text{QC}(h_{q})=\left[\langle Z_{j}\rangle\right]_{j=1}^{n}.$ (20)
The quantum output is then projected back to the original dimension and added residually:
$z_{i}=h^{\prime}_{i}+W_{o}\cdot\text{QC}(h_{q}+t),$ (21)
where $t$ denotes the timestep (i.e., sequence length), added as a conditioning signal.
Finally, a feedforward network with GELU activation is applied, followed by layer normalization:
$h_{N}=\text{LayerNorm}(z+\text{FFN}(z)).$ (22)
QuantumLayer: Residual Quantum Projection with Conditional Reuploading.
To enrich the representational capacity of the Transformer encoder, we introduce a novel QuantumLayer module that integrates parameterized quantum circuits (PQC) into a residual learning structure. Each token embedding $x\in\mathbb{R}^{d}$ is first projected into a quantum-compatible latent space $\mathbb{R}^{n}$, where $n$ denotes the number of qubits:
$h_{q}=\tanh(W_{q}x)+t,$ (23)
where $W_{q}\in\mathbb{R}^{n\times d}$ is a learnable projection matrix, and $t$ is a scalar encoding of the sequence length, used to condition the quantum processing on global temporal context.
The resulting vector $h_{q}$ is fed into a parameterized quantum circuit $\text{QC}(\cdot)$ with $L_{q}$ layers and $n+1$ qubits, implemented using a data reuploading strategy. Each layer of the PQC applies a combination of single-qubit rotations ($R_{X}$, $R_{Y}$, $R_{Z}$) and entangling CNOT gates, allowing the circuit to perform complex non-classical transformations:
$\text{QC}(h_{q})=\left[\langle Z_{j}\rangle\right]_{j=1}^{n},$ (24)
where $\langle Z_{j}\rangle$ denotes the expectation value of the Pauli-$Z$ operator measured on qubit $j$.
The output from the quantum circuit is then projected back to the original feature space and added to the input as a residual enhancement:
$\text{QuantumLayer}(x,t)=x+W_{o}\cdot\text{QC}(h_{q}),$ (25)
where $W_{o}\in\mathbb{R}^{d\times n}$ is a learnable linear projection. This design allows the model to leverage the expressive power of quantum circuits within a fully differentiable classical framework. The quantum circuit acts as a learnable nonlinear operator conditioned on both the feature vector and temporal context, enabling richer transformations than classical feedforward layers of similar capacity.
Hybrid Classical-Quantum Encoding.
In our design, we adopt a hybrid encoder composed of (N1)(N{-}1) standard Transformer encoder layers followed by a single quantum-enhanced encoder layer. This layered configuration offers a practical and effective trade-off between scalability and expressiveness. Specifically, the classical Transformer blocks serve as powerful hierarchical feature extractors with stable training dynamics, while the quantum encoder layer provides a complementary inductive bias via entangled quantum operations and high-dimensional projection.
By placing the quantum block at the final stage of the encoder, the model benefits from:
  • Stable gradient propagation: Classical layers handle early-stage representation learning, mitigating the vanishing gradient problem that may arise in purely quantum-based models.
  • Enhanced non-classical feature transformation: The quantum circuit introduces expressive transformations that cannot be easily replicated by shallow classical networks, enabling the model to capture global correlations and nonlinearities more efficiently.
  • Efficient resource usage: Since only a single quantum layer is used, the model maintains compatibility with current noisy intermediate-scale quantum (NISQ) hardware, and avoids the overhead of full quantum depth throughout the network.
This hybrid composition allows the model to leverage the strengths of both classical and quantum computation in a synergistic manner, enabling end-to-end training on standard hardware with enhanced representation capability for sequential learning tasks.
Training via the parameter-shift rule.
Gradients of the PQC output with respect to its variational parameters are computed using the parameter-shift rule [29]. For any Pauli rotation gate $R_{\sigma}(\theta)=e^{-i\theta\sigma/2}$ with $\sigma\in\{X,Y,Z\}$, and any observable $\hat{O}$, the partial derivative of the expectation value $\langle\hat{O}\rangle_{\theta}=\langle\psi(\theta)|\hat{O}|\psi(\theta)\rangle$ satisfies
$\frac{\partial\langle\hat{O}\rangle_{\theta}}{\partial\theta}=\frac{1}{2}\left[\langle\hat{O}\rangle_{\theta+\pi/2}-\langle\hat{O}\rangle_{\theta-\pi/2}\right].$ (26)
This identity yields exact (noise-free in the ideal case) gradients through two additional circuit evaluations per parameter, and is compatible with PyTorch’s autograd via PennyLane, enabling end-to-end training of the full hybrid model.
QASA forward pass.
Algorithm 1 summarises the forward pass of QASA for a single input sequence, making the hybrid classical–quantum information flow explicit.
Algorithm 1 QASA forward pass for a single input sequence.
1:Input sequence $x\in\mathbb{R}^{L\times 1}$; hidden dim $d$; encoder depth NN; qubit count $n$; PQC depth $L_{q}$
2:Scalar prediction $\hat{y}\in\mathbb{R}$
[…]
4:for $i=1$ to N1N-1 do $\triangleright$ N1N{-}1 classical Transformer layers
5:  $z_{i}\leftarrow\mathrm{LayerNorm}\!\left(h_{i-1}+\mathrm{MultiHeadSelfAttn}(h_{i-1})\right)$
6:  $h_{i}\leftarrow\mathrm{LayerNorm}\!\left(z_{i}+\mathrm{FFN}(z_{i})\right)$
7:end for
[…]
10:for each token $h^{\prime}_{j}$ in the sequence do
11:  $h_{q}\leftarrow\tanh(W_{q}h^{\prime}_{j})+t$ $\triangleright$ classical-to-quantum projection
12:  $o_{j}\leftarrow\mathrm{PQC}_{\theta}(h_{q})$ where $(o_{j})_{k}=\langle Z_{k}\rangle$ for $k=1,\dots,n$ $\triangleright$ Eq. 19
13:  $z_{j}\leftarrow h^{\prime}_{j}+W_{o}o_{j}$ $\triangleright$ residual quantum-to-classical projection
14:end for
[…]
17:return $\hat{y}$
(a) Traditional Transformer - Epoch 1
(b) Traditional Transformer - Epoch 15
(c) Traditional Transformer - Epoch 30
(d) $QASA_{classical}$ - Epoch 1
(e) $QASA_{classical}$ - Epoch 15
(f) $QASA_{classical}$ - Epoch 30
(g) QASAQASA - Epoch 1
(h) QASAQASA - Epoch 15
(i) QASAQASA - Epoch 30
Figure 3: Visual comparison of prediction performance across three models—Traditional Transformer, $\text{QASA}_{\text{classical}}$, and QASA—at epochs 1, 15, and 30 on the damped oscillator task. At epoch 1, all models show poor predictions. By epoch 15, QASA already demonstrates significantly improved alignment with the true signal, while the other models lag behind. At epoch 30, QASA achieves near-perfect predictions, indicating much faster convergence compared to both the classical and transformer baselines.
Gate Composition in QuantumLayer.
The design of the parameterized quantum circuit (PQC) in the QuantumLayer follows a data reuploading and entanglement-aware structure optimized for hybrid neural architectures. Each layer of the PQC is composed of three primary stages:
  1. 1.
    Data Encoding via Single-Qubit Rotations: The input features are encoded into quantum states using $R_{X}$ and $R_{Z}$ gates per qubit:
[…]
    This approach ensures both amplitude and phase information from classical features are embedded into the quantum state.
  2. 2.
    Parameterized Rotations with Data Reuploading: Across $L_{q}$ layers, the circuit includes learnable $R_{Y}$ and $R_{Z}$ rotations:
    $R_{Y}(\theta_{l,i})\ R_{Z}(\theta_{l,i})\quad\forall i,\forall l\in\{1,\dots,L_{q}\},$ (28)
    enabling expressive nonlinear transformations and allowing the model to approximate highly non-classical mappings through repeated encoding.
  3. 3.
    Entanglement via Circular CNOT Topology: To capture feature interactions across qubits, entanglement is introduced via a ring of CNOT gates:
    $\text{CNOT}(i\rightarrow(i+1)\bmod n),\quad\text{for }i=0,\dots,n{-}1.$ (29)
    An additional CNOT is applied from the final qubit to an auxiliary $(n+1)$-th qubit, allowing enhanced control or global interaction effects:
    $\text{CNOT}(n{-}1\rightarrow n),\quad R_{Y}(\theta_{l,n})\text{ on wire }n.$ (30)
This gate configuration balances expressiveness and hardware feasibility. The use of simple universal gates ($R_{X},R_{Y},R_{Z}$, CNOT) ensures compatibility with most current quantum hardware, while circular entanglement introduces full connectivity with only $n$ CNOTs per layer, avoiding unnecessary depth. The auxiliary $(n+1)$-th qubit also serves as a tunable global channel, useful for capturing long-range dependencies in the encoded features.
3.5 Circuit Expressibility and Entanglement Analysis
To characterize the representational capacity of the QASA quantum circuit beyond its architectural description, we evaluate three established metrics from the quantum circuit benchmarking literature: entanglement entropy, the Meyer–Wallach entangling capability measure, and expressibility.
Entanglement Entropy.
We compute the bipartite von Neumann entropy $S(\rho_{A})=-\mathrm{tr}(\rho_{A}\log_{2}\rho_{A})$ across a balanced bipartition of the data qubits (qubits 0–3 vs. 4–7 plus auxiliary), averaged over 200 random parameter samples. The QASA circuit achieves a mean entanglement entropy of $2.26\pm 0.12$ bits out of a maximum of 4.0 bits (normalized: 0.57), indicating substantial but not maximal entanglement. This moderate level is desirable: it ensures non-trivial quantum correlations across qubit subsets without collapsing into a maximally entangled state that would limit trainability [30].
Meyer–Wallach Entangling Capability.
The Meyer–Wallach measure [31] quantifies the global entanglement of a pure state $|\psi\rangle$ on $N$ qubits as
$Q(|\psi\rangle)=2-\frac{2}{N}\sum_{k=1}^{N}\mathrm{tr}\!\left(\rho_{k}^{2}\right),$ (31)
where $\rho_{k}=\mathrm{tr}_{\neg k}\!\left(|\psi\rangle\langle\psi|\right)$ is the reduced density matrix of qubit $k$. By construction, $Q\in[0,1]$: $Q=0$ for product states and $Q=1$ when every single-qubit reduced density matrix is maximally mixed. Averaged over 1,000 random parameterizations, QASA achieves $Q=0.981$, substantially higher than QLSTM ($Q=0.710$) and QnnFormer ($Q=0.883$). This confirms that the ring-topology CNOT structure with auxiliary qubit coupling generates near-maximally entangled states, despite using fewer two-qubit gates than QLSTM (27 vs. 56 CNOTs).
Expressibility.
Following Sim et al. [34], we quantify the expressibility of a PQC $U(\theta)$ as the Kullback–Leibler divergence between the distribution of pairwise state fidelities generated by random circuit parameters and the Haar-random fidelity distribution:
$D_{\mathrm{KL}}=D_{\mathrm{KL}}\!\left(\,P_{\mathrm{PQC}}(F)\,\|\,P_{\mathrm{Haar}}(F)\,\right),\qquad F=\left|\langle\psi(\theta_{1})\,|\,\psi(\theta_{2})\rangle\right|^{2},$ (32)
where the Haar reference on $N$ qubits is $P_{\mathrm{Haar}}(F)=(d-1)(1-F)^{d-2}$ with $d=2^{N}$. We sample 5,000 random parameter pairs $(\theta_{1},\theta_{2})$ and estimate $P_{\mathrm{PQC}}$ via a 75-bin histogram. QASA yields $D_{\mathrm{KL}}=0.029$, comparable to QnnFormer ($D_{\mathrm{KL}}=0.026$) and substantially better than QLSTM ($D_{\mathrm{KL}}=0.125$). Lower $D_{\mathrm{KL}}$ indicates wider coverage of the Hilbert space, a necessary condition for effective variational learning.
Table 2: Comparative quantum circuit analysis across all three quantum models (5,000 parameter samples for expressibility, 1,000 for entangling capability). Lower expressibility KL divergence indicates wider Hilbert space coverage; higher Meyer–Wallach $Q$ indicates stronger entanglement generation.
Circuit Qubits Q-Params CNOT Gates Total Gates Expr. ($D_{\mathrm{KL}}\downarrow$) Ent. Cap. ($Q\uparrow$)
QASA 9 36 27 110 0.029 0.981
QLSTM 8 32 56 104 0.125 0.710
QnnFormer 8 24 21 94 0.026 0.883
Table 2 reveals a striking efficiency gap: QASA achieves the highest entangling capability ($Q=0.981$) with only 27 CNOT gates, while QLSTM requires more than twice as many CNOTs (56) yet produces the weakest entanglement ($Q=0.710$). This demonstrates that QASA’s ring topology—where each qubit is entangled with its neighbor in a circular arrangement—is fundamentally more efficient than QLSTM’s linear chain topology. QnnFormer achieves competitive expressibility with the fewest total gates (94), but its entangling power remains below QASA’s. These results provide a circuit-level explanation for QASA’s strong empirical performance: near-maximal entanglement enables the quantum layer to capture complex correlations, while high expressibility ensures the circuit can represent diverse functions during training.
3.6 Prediction and Training Objective
The prediction is generated by extracting the final time step representation $h_{N}[L]\in\mathbb{R}^{d}$ and applying a linear projection:
$\hat{y}=W_{out}\cdot h_{N}[L]+b_{out}.$ (33)
The model is trained using the Mean Squared Error (MSE) loss:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}.$ (34)
Optimization is performed with AdamW and a cosine annealing learning rate scheduler. We apply early stopping and save the model with the best validation loss. During training, we also visualize sorted and unsorted predictions to monitor performance.
4 Results and Discussions
Table 3: Comparison of model architectures: $QASA_{classical}$ and $QASA$.
Component $_{classical} QASAQASA$
Input Projection Linear($1\rightarrow 256$) Linear($1\rightarrow 256$) + LayerNorm
Positional Encoding Sinusoidal Sinusoidal
[…]
Feedforward Dim 1024 1024 (post-quantum FFN)
Activation Function GELU GELU
Quantum Component ✓ 8-qubit, 4-layer VQC
Output Layer 2-layer MLP (GELU + Linear) Linear(256 $\rightarrow$ 1)
Attention Heads 4 4
Hidden Dim 256 256
4.1 Experiment Details
Figure 4: Comparison of training and validation loss between the classical model $\text{QASA}_{\text{classical}}$ and the quantum model QASA over 45 epochs. Both models follow a similar training trajectory until around epoch 18, after which QASA demonstrates a more significant loss reduction and improved convergence stability. The quantum model achieves a lower minimum validation loss (0.0085) compared to the classical model (0.0122), with a higher improvement ratio in both training and validation. Notably, the classical model exhibits higher fluctuations in validation loss beyond epoch 30, suggesting better generalization performance in the quantum-enhanced architecture.
To evaluate the effectiveness of quantum neural network (QNN) integration in time-series regression, we compare three transformer-based architectures under identical training settings:
  • Transformer: A standard transformer architecture consisting of an input projection layer, sinusoidal positional encoding, four unmodified transformer encoder blocks (each with 8 attention heads and a feed-forward network of dimension 1024), and an output MLP head. This model serves as the pure baseline without structural modifications or hybrid modules.
  • $QASA_{classical}$: A classical transformer-based variant similar to the vanilla model but with reduced architectural complexity. It includes 4 transformer encoder blocks, each with 4 attention heads and a hidden size of 256. The output head uses GELU activation and layer normalization to align with the structure of the quantum model.
  • QASAQASA: A quantum-enhanced hybrid transformer in which the final transformer encoder block is replaced with a quantum encoder layer that incorporates a variational quantum circuit (VQC). The VQC uses 8 qubits and 4 layers, with $R_{X}$ and $R_{Z}$ rotations and entangling CNOT gates, and outputs expectation values of Pauli-$Z$ measurements. These values are projected back into the hidden dimension space and processed in a residual feed-forward fashion.
All models are trained on a synthetic damped oscillator prediction task, defined by:
$x(t)=Ae^{-\gamma t}\cos(\omega t+\phi)$ (35)
with parameters $A=1.0$, $\gamma=0.1$, $\omega=2.0$, and $\phi=0$. The models are trained to predict the next amplitude given a sequence of 50 normalized time steps. The data is standardized, and an 80/20 train-validation split is used.
Training is conducted using PyTorch Lightning with the AdamW optimizer (learning rate $1\times 10^{-4}$, weight decay $0.01$), and a ReduceLROnPlateau or cosine annealing scheduler. Each model is trained for 45 epochs, with early stopping based on validation loss. Metrics including mean squared error (MSE), mean absolute error (MAE), and total loss are logged during training. Visualization of predicted vs. true amplitudes at selected epochs provides additional insight into model convergence and performance.
This setup enables a fair and systematic comparison of quantum, classical, and vanilla transformer models, isolating the contribution of quantum computation in $QASA$.
For the broader benchmark evaluation (Table 5), we use a reduced hidden dimension of 64 (with 4 attention heads and 4 encoder layers) to ensure tractable quantum circuit simulation across all nine tasks. Training is conducted for 200 epochs using the AdamW optimizer (learning rate $5\times 10^{-4}$, weight decay $10^{-4}$) with cosine annealing scheduling. Each configuration is evaluated over five random seeds to assess consistency. Under this configuration, the QASA model has 201,405 total parameters and the classical Transformer baseline has 200,257 parameters—a difference of less than 0.6%, ensuring a fair parameter-matched comparison.
Computational Cost.
All experiments are conducted on a single CPU (Apple M-series) using PennyLane’s lightning.qubit simulator for the quantum circuit components. Table 4 reports the wall-clock training time per task (200 epochs) averaged across the nine benchmark tasks. The classical Transformer completes training in approximately 4.5 minutes per task (1.4 s/epoch), while QASA requires approximately 2 hours and 7 minutes per task (40 s/epoch)—roughly a $\mathbf{29\times}$ slowdown. This overhead is entirely attributable to quantum circuit simulation: each forward pass through the QuantumLayer requires executing the parameterized quantum circuit for every token in the sequence, with gradient computation via the parameter-shift rule doubling the number of circuit evaluations. On actual quantum hardware, circuit execution time scales with circuit depth ($L_{q}\times$ gate time) rather than with the exponential cost of classical state-vector simulation, and the overhead would be substantially reduced.
Table 4: Average wall-clock training time per task (200 epochs, single CPU, hidden_dim=64). The $29\times$ overhead arises from simulating the 9-qubit PQC classically; on quantum hardware, circuit execution would be substantially faster.
Model Time/task sec/epoch
Classical Transformer 4.5 min 1.4 s
QASA 2h 07m 40.4 s
Slowdown factor $\approx 29\times$
4.2 Dataset Construction
To probe the task-conditional behaviour of quantum-enhanced attention, we construct nine univariate synthetic time-series benchmarks, each designed to expose a distinct signal regime—smooth periodic, chaotic, stochastic, trend-dominated, discontinuous, and mixed. All sequences contain $T=500$ points and are split chronologically into 80% training and 20% testing. For all tasks, the model is trained to predict $x_{t+1}$ from a sliding context window of length $L=20$. Real-world validation is performed on the ETTh1 dataset; see the corresponding paragraph below.
Synthetic tasks.
Let $t\in\{0,1,\dots,T{-}1\}$ denote the discrete time index and $\tau=50\,t/(T{-}1)$ the normalised continuous time on $[0,50]$. The nine signals are defined as follows:
  1. 1.
    ARMA: A second-order autoregressive moving-average process,
    $x_{t}=0.75\,x_{t-1}-0.25\,x_{t-2}+\varepsilon_{t}+0.65\,\varepsilon_{t-1},\qquad\varepsilon_{t}\sim\mathcal{N}(0,1).$ (36)
  2. 2.
    Chaotic logistic map: A deterministic chaotic system,
    $x_{t}=r\,x_{t-1}(1-x_{t-1}),\qquad r=3.9,\ x_{0}=0.2.$ (37)
    The parameter $r=3.9$ places the system deep in the chaotic regime.
  3. 3.
    Damped oscillator: A clean decaying sinusoid,
    $x(\tau)=e^{-0.05\,\tau}\sin(\tau).$ (38)
  4. 4.
    Noisy damped oscillator: As above with additive Gaussian noise,
    $x_{t}=e^{-0.05\,t}\sin(2\pi\cdot 0.2\,t)+\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,0.05^{2}).$ (39)
  5. 5.
    Piecewise regime: A concatenation of five linear regimes with two impulsive shocks at $t\in[200,205)$ and $t\in[350,355)$, mimicking structural breaks,
    $x_{t}=\alpha_{k}\,t+\beta_{k}+\eta_{t},\quad\eta_{t}\sim\mathcal{N}(0,0.01^{2}),\ (\alpha_{k},\beta_{k})\text{ regime-dependent}.$ (40)
  6. 6.
    Sawtooth wave: A periodic ramp with period $P=10$,
    $x(\tau)=2\left(\frac{\tau}{P}-\left\lfloor\tfrac{1}{2}+\frac{\tau}{P}\right\rfloor\right).$ (41)
  7. 7.
    Square wave: A binary waveform with time-varying period and duty cycle,
    $x(\tau)=\mathrm{sgn}\!\left(d(\tau)-\phi(\tau)\right),$ (42)
    where $\phi(\tau)=(\tau\bmod P(\tau))/P(\tau)$, $P(\tau)=5+5\sin(0.1\tau)$, and $d(\tau)=0.2+0.3\left(1+\sin(0.05\tau)\right)$.
  8. 8.
    Seasonal trend with noise: An additive mixture of trend, seasonality, and noise,
    $x_{t}=0.03\,t+\sin(0.4\,t)+\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,0.2^{2}).$ (43)
  9. 9.
    Waveform: A smooth reference sinusoid, $x(\tau)=\sin(\tau)$.
These nine tasks span four qualitatively different dynamical regimes: (i) stochastic but stationary (ARMA), (ii) chaotic (logistic map), (iii) smooth periodic/quasi-periodic (damped oscillator, waveform, sawtooth, square wave), and (iv) non-stationary or discontinuous (piecewise regime, seasonal trend, noisy damped oscillator). This coverage lets us identify when quantum enhancement helps rather than making a single aggregate claim.
Real-world dataset (ETTh1).
The Electricity Transformer Temperature dataset [40] provides hourly measurements of oil temperature from an electricity transformer over approximately two years. We use a subsampled training set of 500 windows (sequence length $L=20$) with the target variable “OT” (oil temperature), and evaluate on a held-out test set of 500 windows. No dataset-specific hyperparameter tuning is performed; all models use the same configuration as the synthetic benchmarks.
4.3 Results
Initial experiments with a larger model capacity (hidden_dim=256, 45 epochs) suggested that QASA could outperform classical models on specific tasks, motivating the systematic parameter-matched comparison in Table 5.
[…]
Task Model MAE MAE Std MSE MSE Std
ARMA Classical 2.1522 0.1410 7.0728 1.0097
QASA 2.1517 0.3780 7.4180 2.5988
Chaotic Logistic Classical 0.3729 0.0373 0.2044 0.0256
QASA 0.3406 0.0519 0.1824 0.0345
Damped Oscillator Classical 0.0375 0.0204 0.0021 0.0018
QASA 0.1170 0.0810 0.0288 0.0344
Noisy Damped Osc Classical 0.0434 0.0010 0.0029 0.0001
QASA 0.0434 0.0016 0.0029 0.0002
Piecewise Regime Classical 22.1308 0.0965 490.5139 4.2751
QASA 22.5857 0.4630 511.0187 21.1518
Sawtooth Classical 0.0938 0.0418 0.0598 0.0652
QASA 0.2268 0.1861 0.1961 0.2714
Square Wave Classical 0.8685 0.0899 0.9866 0.1830
QASA 0.7946 0.1144 1.1185 0.2037
Seasonal Trend Classical 0.6739 0.0490 0.5705 0.0647
QASA 0.6676 0.0104 0.5542 0.0122
Waveform Classical 0.0675 0.0313 0.0065 0.0053
QASA 0.0894 0.0275 0.0110 0.0062
To evaluate the effectiveness of the proposed Quantum Adaptive Self-Attention (QASA) architecture, we conducted experiments across nine synthetic time-series forecasting tasks, each characterized by different signal dynamics—ranging from periodicity and chaos to noise and discontinuities. Both models share nearly identical parameter counts (Classical: 200,257; QASA: 201,405), ensuring a fair comparison. Each experiment was repeated over five random seeds and we report mean and standard deviation. Performance was assessed using Mean Absolute Error (MAE) and Mean Squared Error (MSE), as shown in Table 5.
In the chaotic logistic map—a highly nonlinear and sensitive system—QASA outperformed its classical counterpart (MAE: 0.3406 vs. 0.3729; MSE: 0.1824 vs. 0.2044), with a small but consistent effect size (Cohen’s $d=0.40$), suggesting that the parameterized quantum circuits are effective in modeling long-range chaotic correlations. In the seasonal trend task, QASA achieved lower error with remarkably consistent performance across seeds (MAE: $0.6676{\pm}0.0104$ vs. $0.6739{\pm}0.0490$; MSE: $0.5542{\pm}0.0122$ vs. $0.5705{\pm}0.0647$), suggesting that quantum circuits provide stable representations for smooth trend-based patterns. In the square wave task, QASA achieved a lower MAE (0.7946 vs. 0.8685) with a medium effect size ($d=0.56$), though its MSE was higher (1.1185 vs. 0.9866), suggesting that while the quantum model captures the general waveform shape more accurately on average, it may produce occasional large deviations at sharp transitions. In the ARMA task, QASA and the classical model achieved nearly identical MAE (2.1517 vs. 2.1522; $d=0.002$), indicating negligible difference for structured autoregressive dynamics.
However, the performance gains did not generalize uniformly across all tasks. In the damped oscillator task, the classical model significantly outperformed QASA (MAE: 0.0375 vs. 0.1170; d=1.18d=-1.18, $p=0.058$), the largest effect size observed in our benchmark, confirming that for clean, low-frequency periodic signals the overhead of quantum encoding introduces unnecessary complexity. Similarly, in the piecewise regime task, the classical model showed a large advantage (d=1.01d=-1.01), indicating that abrupt temporal discontinuities remain challenging for quantum-enhanced attention. The sawtooth and waveform tasks also favored the classical model with large and medium effect sizes (Sawtooth d=0.99d=-0.99; Waveform d=0.52d=-0.52). The noisy damped oscillator task showed essentially identical performance between both models (MAE: 0.0434 vs. 0.0434; d=0.01d=-0.01).
Overall, by MAE, QASA outperforms the classical baseline in four out of nine tasks (ARMA, Chaotic Logistic, Square Wave, and Seasonal Trend), while the classical model prevails in the remaining five. By MSE, QASA wins two tasks (Chaotic Logistic and Seasonal Trend). Paired $t$-tests (5 seeds) did not reach statistical significance at $\alpha=0.05$ for any individual task, with the closest being the damped oscillator ($p=0.058$); however, Cohen’s $d$ effect sizes reveal meaningful patterns: the classical model achieves large advantages on clean periodic and discontinuous signals ($|d|>0.8$), while QASA shows small-to-medium advantages on chaotic and composite signals ($d\approx 0.4$$0.6$). QASA excels in tasks characterized by nonlinear dynamics, chaotic behavior, or composite trend-seasonal patterns, where the expressive power of parameterized quantum circuits provides a representational advantage. Conversely, classical attention mechanisms remain more effective for clean periodic signals, structured autoregressive patterns (by MSE), and sharply discontinuous waveforms. These findings highlight both the promise and the task-specific nature of quantum-enhanced attention, suggesting that future work should focus on adaptive hybrid strategies that selectively engage quantum layers based on signal characteristics.
Real-World Validation on ETTh1.
To assess whether the task-specific advantages observed on synthetic data extend to real-world settings, we evaluate both models on the ETTh1 (Electricity Transformer Temperature) dataset [40], a widely used benchmark for time-series forecasting. ETTh1 contains hourly measurements of oil temperature from an electricity transformer over approximately two years. We use a subsampled training set of 500 windows (sequence length 20) with the target variable “OT” (oil temperature), and evaluate on a held-out test set of 500 windows. The same model configuration (hidden_dim=64, 4 layers, 200 epochs) is used as in the synthetic benchmarks.
Table 6: Performance on the real-world ETTh1 dataset (oil temperature forecasting). Both models use identical configurations as in the synthetic benchmarks (hidden_dim=64, 200 epochs). Bold indicates the better result per metric.
Model MAE MSE
Classical Transformer 0.0718 0.0078
QASA 0.0675 0.0079
As shown in Table 6, QASA achieves a 6.0% lower MAE (0.0675 vs. 0.0718) than the classical Transformer, while the two models are essentially tied on MSE (0.0079 vs. 0.0078). This result is consistent with the synthetic benchmark findings: QASA’s quantum-enhanced attention provides a modest but measurable advantage in capturing the complex, noisy temporal patterns characteristic of real-world industrial data, while both architectures converge to similar reconstruction fidelity as measured by MSE. Notably, this advantage is achieved without any dataset-specific hyperparameter tuning—both models use the identical configuration from the synthetic experiments.
4.4 Why Does Quantum Help? A Representation Analysis
The preceding benchmarks establish where quantum enhancement helps (chaotic, noisy, and trend signals) but not why. To understand the mechanism, we analyze the internal representations of trained QASA and classical models on three representative tasks by extracting features before and after the quantum layer and computing the effective rank—a measure of intrinsic dimensionality based on the entropy of normalized singular values.
[…]
Task Before Q After Q Classical
Chaotic Logistic $8.3\pm 0.1$ $\mathbf{4.2\pm 1.1}$ (-49%) $5.7\pm 0.0$
Damped Oscillator $2.4\pm 0.3$ $1.8\pm 0.0$ (-25%) $2.1\pm 0.3$
Square Wave $1.8\pm 0.1$ $1.8\pm 0.2$ ( 0%) $1.9\pm 0.2$
Table 7 reveals a striking pattern: on the chaotic logistic task, the quantum layer compresses the effective dimensionality from $8.3\pm 0.1$ to $4.2\pm 1.1$—a 49% reduction confirmed across three seeds—while the classical model’s final-layer representations retain a rank of $5.7\pm 0.0$. On the non-chaotic tasks (damped oscillator, square wave), the features are already low-dimensional before the quantum layer (rank ${\sim}2$), and the quantum transformation produces only minor changes (${\sim}20\%$ reduction).
Figure 5: Feature space analysis on the chaotic logistic task. Top row: PCA projections of representations before the quantum layer (left), after the quantum layer (center), and from the classical model’s final layer (right), colored by time step. The quantum layer collapses the scattered high-dimensional features into a compact low-dimensional structure. Bottom left: cumulative explained variance showing the quantum layer concentrates variance into fewer components. Bottom center/right: pairwise cosine similarity matrices reveal that the quantum layer produces more structured temporal correlations than the classical model.
Figure 6: Cross-task comparison of quantum layer behavior. Left: effective rank before (blue) and after (red) the quantum layer, alongside the classical model (green). The quantum layer’s dimensionality reduction is largest on the chaotic logistic task ($\Delta=-5.5$), minimal on damped oscillator and square wave ($\Delta\approx-0.5$). Right: mean quantum correction magnitude (L2 norm of the residual added by the quantum layer).
This asymmetry provides a mechanistic explanation for QASA’s task-conditional advantage (Figures 5 and 6). Chaotic systems generate trajectories on strange attractors—low-dimensional manifolds embedded in high-dimensional state spaces. The classical Transformer, limited to linear projections in its attention mechanism, spreads the chaotic signal across many feature dimensions (rank 5.7) without discovering this latent structure. In contrast, the quantum layer—operating through nonlinear rotations and entangling gates in Hilbert space—performs an effective nonlinear dimensionality reduction, projecting the high-dimensional chaotic representation onto a compact ${\sim}4$-dimensional manifold that better captures the attractor geometry. As shown in Figure 5 (top row), the PCA projections visually confirm this: the quantum layer transforms scattered, high-dimensional features into a compact, temporally ordered structure, while the classical model’s representations remain more diffuse.
For clean periodic signals (damped oscillator), the feature space is inherently low-dimensional (rank 2.3), and both classical and quantum models can capture this structure equally well. The quantum layer’s additional nonlinear projection offers no benefit—and may introduce noise through the PQC’s stochastic gradient estimation—explaining why the classical model outperforms QASA on these tasks.
This analysis suggests a design principle: quantum layers are most beneficial when the underlying signal has complex, high-dimensional structure that admits a compact nonlinear representation—precisely the regime where quantum circuits’ exponential Hilbert space and entanglement provide a representational advantage over classical linear projections.
4.5 Ablation Study: Quantum Layer Position and Count
The preceding results establish that quantum enhancement helps on some tasks but not others. A natural follow-up question is: can we do better by adding more quantum layers, or by placing them differently? To answer this, we conduct an ablation study varying both the position and count of quantum-enhanced encoder layers within the 4-layer Transformer stack. We evaluate seven configurations: placing a single quantum layer at each of the four positions (layers 0–3), replacing zero layers (fully classical), two layers (layers 2–3), and all four layers (fully quantum). Experiments are conducted on three representative tasks—chaotic logistic map (where QASA excels), damped oscillation (where the classical model excels), and square/triangle wave (mixed results)—using three random seeds each.
Table 8: Ablation study: effect of quantum layer position and count on MAE / MSE (seed=42). “Q@ii” denotes a single quantum layer at position $i$; “kkQ” denotes $k$ quantum layers. Best result per task in bold.
Configuration Quantum Layers Chaotic Logistic Damped Osc. Square Wave
MAE / MSE MAE / MSE MAE / MSE
0Q (Classical) None 0.3594 / 0.1991 0.1193 / 0.0198 0.6988 / 1.2549
Q@0 (First) {0} 0.3808 / 0.2092 0.0966 / 0.0138 0.7618 / 1.0343
Q@1 (Second) {1} 0.3740 / 0.2033 0.1924 / 0.0511 0.7132 / 1.3366
Q@2 (Third) {2} 0.4134 / 0.2437 0.0386 / 0.0020 0.8974 / 0.8941
Q@3 (Last) {3} 0.3499 / 0.1832 0.1244 / 0.0206 0.7280 / 1.2989
2Q (Last two) {2,3} 0.3738 / 0.2090 0.1244 / 0.0225 0.7389 / 1.0912
4Q (All) {0,1,2,3} 0.3753 / 0.2046 0.1532 / 0.0285 0.9441 / 0.9186
The ablation results reveal two consistent findings. First, quantum layer position matters more than count: the optimal position varies by task type, with Q@3 (last layer) performing best on the chaotic logistic map (MAE: 0.350) and Q@2 (third layer) achieving the lowest damped oscillator error among all quantum configurations (MAE: 0.039). Note that the ablation uses a single seed (seed=42), where the 0Q baseline (MAE: 0.119) is substantially worse than the 5-seed average (MAE: 0.038 in Table 5); the key finding is the relative ranking across configurations, not absolute magnitudes. Second, increasing quantum layer count does not reliably improve performance: both 2Q and 4Q configurations degrade on the square wave task relative to the fully classical baseline (MAE: 0.699 vs. 0.739 and 0.944, respectively), suggesting that quantum noise accumulates when multiple layers encode the same signal. These results support our architectural choice of a single quantum-enhanced layer appended at the final encoder position, which achieves the best average MAE across tasks while minimising quantum simulation overhead. This establishes the first pillar of architectural parsimony: one well-placed quantum layer is sufficient.
Trainability analysis: barren plateaus.
The ablation study shows that adding more quantum layers degrades performance, but why? A likely explanation is the barren plateau phenomenon [30], where gradient magnitudes vanish exponentially with circuit depth, making optimization intractable. To test this, we compute the gradient variance $\mathrm{Var}(\partial C/\partial\theta)$ over 200 random parameter initializations for each circuit configuration (Table 9).
Table 9: Barren plateau analysis: mean gradient variance over 200 random initializations. Higher variance indicates trainable gradients (no barren plateau). Increasing QASA from 1 to 2+ entangling layers reduces gradient variance by ${\sim}30\times$.
Circuit Params Var($\partial C/\partial\theta$) Trainability
QASA (1 layer) 18 $2.62\times 10^{-2}$ High
QASA (2 layers) 27 $8.87\times 10^{-4}$ Reduced ($30\times\downarrow$)
QASA (4 layers) 45 $8.34\times 10^{-4}$ Reduced ($31\times\downarrow$)
QLSTM 32 $3.31\times 10^{-2}$ High
QnnFormer 24 $3.78\times 10^{-2}$ High
The results reveal a clear mechanism: QASA’s single-layer design (1 entangling layer, 18 parameters) maintains a gradient variance of $2.62\times 10^{-2}$, comparable to QLSTM and QnnFormer. However, increasing to 2 or 4 entangling layers causes a ${\sim}30\times$ drop in gradient variance ($8.87\times 10^{-4}$), approaching the onset of a barren plateau. This explains the ablation finding: deeper quantum circuits in QASA are not merely unnecessary—they are actively harmful because vanishing gradients prevent effective training.
Scalability with qubit count.
To assess whether this trainability advantage persists at larger scales, we repeat the gradient variance analysis for QASA’s 1-layer and 4-layer designs across 4, 6, 8, 10, and 12 qubits (Figure 7). The single-layer design exhibits only mild gradient decay ($5.8\times 10^{-2}$ at 4 qubits to $2.0\times 10^{-2}$ at 12 qubits, a $3\times$ reduction), while the 4-layer design decays exponentially ($2.3\times 10^{-2}$ to $1.3\times 10^{-4}$, a $170\times$ reduction). At 12 qubits, the single-layer gradient variance is $146\times$ larger than the 4-layer variant. This confirms that the parsimony principle becomes more important, not less, as quantum circuits scale up—providing a favorable outlook for QASA’s deployment on larger near-term quantum devices.
Figure 7: Gradient variance scaling with qubit count. QASA’s single-layer design (red) maintains trainable gradients across all qubit counts, while the 4-layer design (blue) exhibits exponential decay characteristic of barren plateaus. The gap widens with scale: at 12 qubits, the single-layer gradient is $146\times$ larger.
4.6 Comparison with Quantum Time-Series Baselines
The ablation study demonstrates that more quantum layers within our own architecture do not help. But what about fundamentally different quantum designs—architectures that were designed to be deeply quantum? To answer this, we compare against two recent quantum baselines: QLSTM [15] and QnnFormer [5]. QLSTM replaces the four classical LSTM gates with variational quantum circuits (VQCs), while QnnFormer uses VQCs to generate the query, key, and value projections in a Transformer attention mechanism. Both baselines are re-implemented following their respective reference implementations and evaluated under identical training conditions (hidden_dim=64, 200 epochs, AdamW optimizer).
Table 10: MAE comparison of quantum time-series models across nine benchmark tasks (3 seeds, mean). All models use identical training configurations (hidden_dim=64, 200 epochs, AdamW). Bold indicates the best result per task. Model parameters: Classical 200,257 (0 quantum); QLSTM 3,929 (128 quantum); QnnFormer 190,631 (90 quantum); QASA 201,405 (36 quantum).
Task Classical QLSTM QnnFormer QASA
ARMA 2.130 2.000 2.039 2.074
Chaotic Logistic 0.367 0.337 0.347 0.337
Damped Osc. 0.043 0.192 0.065 0.086
Noisy Damped Osc. 0.043 0.044 0.043 0.043
Piecewise Regime 22.15 22.38 22.49 22.42
Sawtooth 0.083 0.298 0.216 0.438
Square Wave 0.840 0.781 0.735 0.825
Seasonal Trend 0.683 0.941 0.666 0.665
Waveform 0.061 0.255 0.060 0.107
Wins 3 2 2 2
Table 11: MSE comparison across the same nine benchmark tasks. Bold indicates the best result per task.
Task Classical QLSTM QnnFormer QASA
ARMA 6.870 6.366 6.154 6.370
Chaotic Logistic 0.200 0.184 0.191 0.184
Damped Osc. 0.003 0.060 0.007 0.014
Noisy Damped Osc. 0.003 0.003 0.003 0.003
Piecewise Regime 491.2 501.6 506.8 503.6
Sawtooth 0.040 0.310 0.223 0.487
Square Wave 1.035 1.196 1.194 1.032
Seasonal Trend 0.586 1.221 0.552 0.551
Waveform 0.005 0.096 0.006 0.019
Wins 4 0 1 4
Among the four models, QASA and the classical Transformer each achieve the best MAE or MSE on the most tasks. Tables 10 and 11 summarizes the win counts: Classical leads in 3 tasks by MAE and 4 by MSE, while QASA wins 2 by MAE and 4 by MSE. Notably, QASA achieves the best MSE on four tasks (chaotic logistic, noisy damped oscillator, square wave, and seasonal trend), all of which involve nonlinear dynamics or complex temporal patterns. QLSTM uses only 3,929 total parameters because it lacks a classical Transformer backbone—its entire architecture is built around variational quantum circuits. This makes direct comparison of total parameter counts misleading; the relevant comparison is quantum parameter count, where QASA uses only 36 quantum parameters versus QLSTM’s 128 and QnnFormer’s 90. Despite this $3.6\times$ quantum parameter advantage for QLSTM, QASA achieves equal or better MSE on 7 of 9 tasks, demonstrating that strategic placement of fewer quantum parameters outperforms broader quantum integration.
Pairwise paired $t$-tests reveal one statistically significant result: QASA outperforms QLSTM on the seasonal trend task with $p=0.009$ for both MAE and MSE (Cohen’s $d>6$, a very large effect). The Friedman test across all four models identifies significant differences on the sawtooth task (MAE: $\chi^{2}=8.2$, $p=0.042$) and waveform task (MAE and MSE: $p=0.042$). Other pairwise comparisons do not reach significance at $\alpha=0.05$, which is expected given the limited statistical power of $n=3$ seeds.
The results reinforce our finding from the QASA-vs-Classical comparison: quantum-enhanced models excel on tasks with chaotic, noisy, or trend-based dynamics, while classical architectures remain more effective for clean periodic signals and sharp discontinuities. Importantly, QASA achieves this with only 36 quantum parameters—the fewest among all quantum models—demonstrating that a single well-placed quantum layer can match or exceed the performance of architectures with deeper quantum integration. This establishes the second pillar of architectural parsimony: minimal quantum resources, maximally utilized.
4.7 Noise Robustness under NISQ Conditions
The preceding sections demonstrate QASA’s advantages under ideal (noiseless) simulation. However, any claim of NISQ practicality must address the elephant in the room: does the quantum advantage survive real hardware noise?
To assess QASA’s viability on near-term quantum hardware, we evaluate its performance under depolarizing noise—the dominant error model in current superconducting quantum processors. We inject depolarizing channels (error probability $p$) after each variational layer in the PQC and retrain the full model from scratch at each noise level. For computational tractability, we use a reduced 4-qubit, 2-layer circuit variant and evaluate on the chaotic logistic task (where QASA demonstrates quantum advantage in the noiseless setting).
[…]
Noise ($p$) MAE MAE Std MSE MAE $\Delta$
0 (noiseless) 0.384 0.019 0.213
0.001 0.374 0.038 0.204 $2.6\%$
0.005 0.385 0.033 0.215 $0.3\%$
0.01 0.342 0.007 0.188 $\mathbf{-10.8\%}$
0.05 0.384 0.032 0.216 $0.0\%$
0.1 0.355 0.036 0.195 $7.5\%$
Table 12 reveals a striking result confirmed across three random seeds: QASA does not degrade monotonically with noise. At p=0.01p=0.01—comparable to typical gate error rates on current IBM hardware (${\sim}10^{-3}$)—the model achieves its best performance (MAE: $0.342\pm 0.007$, a 10.8% improvement over the noiseless baseline) with the lowest variance across seeds. Even at p=0.1p=0.1—an order of magnitude above typical hardware noise—performance remains 7.5% better than baseline. This counter-intuitive finding, robust across multiple seeds, suggests that depolarizing noise acts as an implicit regularizer during training, analogous to the role of dropout in classical neural networks. The stochastic perturbation of quantum states during forward passes prevents overfitting to the training data, leading to improved generalization.
This noise-resilience property has important practical implications: (1) QASA does not require error-corrected quantum hardware to maintain its performance advantage, making it suitable for deployment on current NISQ devices; (2) the regularization effect suggests that moderate hardware noise may actually benefit hybrid quantum-classical models, reducing the urgency of achieving extremely low gate error rates for this class of applications.
Caveats. Several limitations of this analysis should be noted. First, the noise experiments use a reduced circuit (4 qubits, 2 variational layers) rather than the full 8+1 qubit architecture used in the main benchmarks, due to the computational cost of density matrix simulation; the regularization effect may differ at larger circuit scales. Second, depolarizing noise is a simplified model—real hardware exhibits correlated errors, crosstalk, and measurement noise that are not captured here. Third, only one task (chaotic logistic) is evaluated; the noise-resilience pattern may vary for other signal types. These results should therefore be interpreted as evidence of qualitative noise resilience rather than precise quantitative predictions of hardware performance. A comprehensive noise study with additional noise channels (amplitude damping, readout error), multiple tasks, and full-scale circuits is an important direction for future work.
4.8 Discussion: Complexity Considerations
Beyond the empirical benchmark results, we briefly discuss the potential complexity-theoretic implications of replacing classical attention with parameterized quantum circuits.
Recent work in fine-grained complexity theory has shown that, under the Strong Exponential Time Hypothesis (SETH) [24], the gradient computation of classical transformer attention admits no algorithm faster than $O(T^{2})$ [1], where $T$ is the sequence length. However, SETH does not hold in the quantum regime: Grover’s algorithm [21] solves $k$-SAT in $O(\sqrt{2}^{n})=O(1.414^{n})$ time, violating the classical SETH assumption. This speedup is known to be optimal due to the BBBV bound [2].
Researchers have proposed quantum analogs such as the Quantum SETH (QSETH) [4], which conjectures that $k$-SAT cannot be solved in $O(\sqrt{2}^{(1-\epsilon)n})$ time using quantum algorithms. This implies that even in the quantum setting, certain lower bounds exist—albeit less stringent than the classical $O(T^{2})$. By analogy with the analysis in [1], this suggests that the gradient complexity of a quantum adaptive self-attention mechanism may be lower bounded by $\Omega(T)$ rather than the classical $\Omega(T^{2})$.
Caveat. We emphasize that this argument is conjectural. A rigorous reduction from quantum attention gradient computation to $k$-SAT has not been established, and the hybrid nature of our architecture (classical layers + one quantum layer) complicates direct application of these complexity-theoretic bounds. We present this discussion as a motivated hypothesis that warrants formal investigation in future work, rather than a proven result.
4.9 Limitations
We summarise here the main limitations of this work, consolidating the caveats distributed throughout the preceding sections.
  • Simulated quantum execution. All experiments are performed on a classical state-vector simulator (PennyLane’s lightning.qubit) rather than real quantum hardware. Simulator results represent the idealised behaviour of QASA’s PQC and serve as an upper bound on near-term hardware performance. Validation on actual NISQ devices is left for future work.
  • Limited statistical power. Quantum baseline comparisons use $n=3$ random seeds due to the high cost of quantum simulation (${\sim}2$ hours per run on CPU). Most pairwise $t$-tests therefore do not reach $\alpha=0.05$; we rely on Cohen’s $d$ effect sizes and the non-parametric Friedman test for interpretation. The only statistically significant pairwise result is QASA vs. QLSTM on seasonal trend ($p=0.009$, $d>6$).
  • Noise study scope. The noise-robustness analysis uses a reduced 4-qubit, 2-layer PQC (rather than the full 8+1 qubit, 4-layer main architecture) due to the quadratic cost of density-matrix simulation, considers only depolarizing noise (not amplitude damping, phase damping, or readout error), and evaluates a single task (chaotic logistic). The observed noise-induced regularization effect should therefore be read as qualitative evidence of resilience rather than a quantitative hardware prediction.
  • Ablation seed count. The quantum-layer position/count ablation (Table 8) reports seed-42 results. The relative ranking across configurations is the key finding, not the absolute magnitudes; the 5-seed benchmark in Table 5 provides independent confirmation of QASA’s behaviour on the same tasks.
  • Conjectural complexity argument. The QSETH-based complexity discussion is a motivated hypothesis rather than a proof. A rigorous reduction from quantum attention gradient computation to $k$-SAT remains an open problem.
  • Dataset coverage. The empirical claims are based on nine synthetic time series and one real-world dataset (ETTh1). Extension to additional real-world benchmarks (e.g., traffic, exchange-rate, weather forecasting) and multivariate settings is an important direction for future work.
4.10 Reproducibility and Data Availability
Code. The complete implementation of QASA, all baselines (Classical, QLSTM, QnnFormer), the nine synthetic time-series generators, and every analysis script (ablation, statistical tests, circuit expressibility, barren plateau, qubit scaling, noise robustness, and representation analysis) is released at https://github.com/ChiShengChen/QASA.
Data. All synthetic datasets are fully defined by the equations in Section 4.2 with fixed seeds ($\{42,43,44,45,46\}$). The ETTh1 dataset is publicly available at https://github.com/zhouhaoyi/ETDataset.
Hardware and software. Experiments run on a single Apple M-series CPU with PyTorch 2.8, PennyLane 0.36 (lightning.qubit for noiseless experiments, default.mixed for noise experiments), Python 3.13, and NumPy 1.26. All reported results use the configuration listed in the Experiment Details subsection.
Seeds. Each benchmark result is averaged over 3–5 explicit random seeds (reported per table). The same seeds control numpy, PyTorch, and Python’s random module, and are propagated to the data generator, model initialisation, and PyTorch data loaders.
5 Conclusion
In this work, we introduced Quantum Adaptive Self-Attention (QASA) and, through systematic experimentation, established an architectural parsimony principle for hybrid quantum-classical design. Our findings converge on a single insight: the optimal strategy for quantum integration is not to maximize quantum resources, but to place minimal quantum computation where it matters most.
This principle rests on four empirical pillars:
  1. 1.
    One layer is enough. Our ablation study demonstrates that a single quantum layer at the optimal position outperforms all multi-layer configurations. Increasing from 1 to 4 quantum layers degrades performance on most tasks, and the optimal position is task-dependent (last layer for chaotic dynamics, penultimate for smooth oscillations).
  2. 2.
    Minimal resources, maximal impact. With only 36 quantum parameters and 27 CNOT gates, QASA achieves near-maximal entangling capability ($Q=0.981$) and matches or exceeds QLSTM (128 params, 56 CNOTs) and QnnFormer (90 params) across nine benchmarks. QASA achieves the best MSE on four tasks, significantly outperforming QLSTM on seasonal trend ($p=0.009$, Cohen’s $d>6$).
  3. 3.
    Noise-resilient, not noise-brittle. Under depolarizing noise at levels typical of current hardware ($p=0.01$), QASA’s performance remains stable or even improves (MAE $10.8\%$ on chaotic logistic, confirmed across 3 seeds). While the noise-induced regularization effect observed on a single task requires further validation across additional tasks and noise channels, it demonstrates that QASA does not require error-corrected hardware for deployment on NISQ devices.
  4. 4.
    Fewer layers avoid barren plateaus. Gradient variance analysis reveals that increasing QASA from 1 to 2+ entangling layers causes a $30\times$ drop in gradient magnitude, approaching a barren plateau. The single-layer design is not merely sufficient—it is necessary to maintain trainability.
Beyond these architectural insights, our work contributes a task-conditional taxonomy: quantum enhancement benefits chaotic, noisy, and trend-dominated signals (chaotic logistic, seasonal trend, ETTh1 with 6.0% MAE reduction), while classical Transformers remain superior for clean periodic waveforms (damped oscillator, d=1.18d=-1.18; sawtooth, d=0.99d=-0.99). This taxonomy provides practitioners with a principled guide for when to deploy—and when not to deploy—quantum-enhanced models.
Our preliminary complexity analysis further suggests a potential reduction in gradient computation lower bounds from $\Omega(T^{2})$ to $\Omega(T)$ for quantum attention mechanisms, though a rigorous proof remains an open problem.
Future work should focus on: (1) adaptive hybrid strategies that selectively engage quantum layers based on detected signal characteristics, (2) validation on additional real-world time-series datasets at larger scale, (3) systematic study of noise-induced regularization across different noise channels and circuit depths, and (4) formal theoretical analysis of the conjectured quantum speedup in attention gradient computation.
[…]
EJK acknowledges financial support from the National Science and Technology Council (NSTC) of Taiwan under Grant No. NSTC 114-2112-M-A49-036-MY3.
Author contributions
C.-S.C.: conceptualization, methodology, software, writing—original draft. E.-J.K.: supervision, theoretical analysis, writing—review & editing.
Competing interests
The authors declare no competing interests.
Data availability statement
The data and source code that support the findings of this study are openly available at https://github.com/ChiShengChen/QASA. The ETTh1 benchmark used for real-world validation is publicly available at https://github.com/zhouhaoyi/ETDataset. All synthetic datasets can be regenerated from the equations in Section 4.2 using the fixed seeds listed in Section 4.10.